# PrivyWave: Privacy-Aware Wireless Sensing of Heart Beat

### v1

@@ 1. Introduction
Wireless sensing technologies have advanced rapidly in recent years. Systems using acoustic (Nandakumar et al., 2015), WiFi (Abdelnasser et al., 2015), and millimeter-wave (mmWave) (Yang et al., 2017) sensors can now monitor vital signs without any physical contact, detecting chest movements as small as few millimeters to measure breathing and heartbeat. These capabilities enable critical healthcare applications, such as detecting sleep apnea (Nandakumar et al., 2015), identifying opioid overdoses (Nandakumar et al., 2019), even detecting cardiac arrest and calling emergency services (Chan et al., 2019) all without requiring users to wear sensors.
However, these same technological capabilities that promise healthcare benefits also create privacy risks. Any ambient device such as a WiFi router, smart speaker, home assistant or someone’s smartphone, can potentially become a covert physiological signal monitor, silently extracting vital signs without user awareness or consent. This threat is particularly concerning because vital signs reveal far more than basic metrics, they could expose emotional states (Zhao et al., 2016), or stress levels (Ha et al., 2021) that may even indicate mental health conditions. In addition, some wireless signals are invisible and penetrate walls and clothing, creating an omnipresent surveillance risk which is hard for users to detect or avoid. So as wireless sensing becomes ubiquitous, developing solutions that maintain its practical benefits while preventing unauthorized surveillance is a necessity.
Existing privacy protection approaches for wireless signals fall into two categories: post-collection data processing and real-time protection. Post-collection approaches (Shi et al., 2010; Hassan et al., 2019; Liu et al., 2023) protect data after acquisition through techniques such as data aggregation (Shi et al., 2010), differential privacy (Hassan et al., 2019), and signal tampering (Liu et al., 2023). However, these methods cannot prevent unauthorized sensing in the first place, as adversaries can collect raw signals in real-time before any post-collection protection is applied. For real-time protection, jamming methods (Chen et al., 2020; Yu et al., 2025; Qiao et al., 2025) inject noise to degrade signal quality, but they are modality-specific—protecting against one type of sensor (e.g., acoustic) but not others (e.g., radio-frequency or RF). Anti-Sensing (Oshim et al., 2025) uses wearable oscillators to mislead radar-based heartbeat detection, but blocks all sensors indiscriminately. VitalHide (Gao et al., 2025) presented the first approach for selective protection using vibration-based obfuscation. However, this work remained conceptual, lacking theoretical foundations, formal privacy analysis, and systematic validation.
@@ 2. Related Work
PriSense (Shi et al., 2010) Sensor Data N/A
Diff. Privacy (Hassan et al., 2019) Sensor Data N/A
mmFilter (Liu et al., 2023) mmWave Radar N/A
@@
Wearable Jammer (Chen et al., 2020) Audio
Dynamic Jamming (Yu et al., 2025) Audio
NUSGuard (Qiao et al., 2025) Voice
@@
RF-Protect (Shenoy et al., 2022) Human Tracking
Radar Obfus. (Argyriou, 2023) Activity
Anti-Sensing (Oshim et al., 2025) Vital Signs
VitalHide (Gao et al., 2025) Vital Signs
@@ 2.1. Wireless Vital Sign Sensing and Privacy Implications
Wireless vital sign sensing has evolved into reliable systems with high accuracy. Many smart devices have gained the ability to accurately detect vital signs without physical contact. Nandakumar et al. (Nandakumar et al., 2015) demonstrated that smartphones can be turned into vital sign monitors by leveraging the acoustic sensors in the device. WiFi-based approaches extract breathing and heartbeat from the RF signals reflected by the subject (Abdelnasser et al., 2015; Liu et al., 2015). The mmWave sensors, commonly deployed for presence detection, can also measure breathing and heartbeat with high precision (Yang et al., 2017; Gong et al., 2021).These sensing capabilities have enabled beneficial applications in everyday life. Researchers have developed systems for sleep apnea detection (Nandakumar et al., 2015), which can alert users to potentially dangerous breathing interruptions during sleep. Contactless vital sign monitoring has been proposed for overdose detection (Nandakumar et al., 2019), enabling earlier intervention in emergency situations. Cardiac arrest detection systems (Chan et al., 2019) can automatically alert emergency services when abnormal vital signs are detected. However, vital signals carry far more information than just physiological measurements. Stress levels can be inferred from vital signals (Ha et al., 2021), emotional states can be detected from combined respiratory and cardiac patterns (Zhao et al., 2016), and cognitive load during mental tasks can be assessed through cardiovascular responses (Solhjoo et al., 2019). As sensing accuracy continues to improve, the precision of these inferences also increases, making the privacy risks more severe.
@@ 2.2.1. Post-Collection Data Processing
There are a set of algorithms that focuses on protecting data after it has been collected by sensors. These approaches assume that sensing has already occurred and apply privacy-preserving techniques during subsequent processing, transmission, or storage stages. PriSense (Shi et al., 2010) introduced privacy-preserving data aggregation for sensor networks using data slicing, where each sensor splits its reading into multiple shares distributed to randomly selected cover nodes, protecting individual data unless the aggregation server colludes with all cover nodes. Differential privacy approaches (Hassan et al., 2019) provide stronger theoretical guarantees by adding calibrated noise to sensor data datasets, ensuring that the presence or absence of any individual’s data cannot be reliably determined while preserving aggregate statistics. For mmWave radar-based sensing, mmFilter (Liu et al., 2023) applies signal reversion techniques that tamper with radar data after collection but before transmission to sensing processors. While these post-collection methods provide valuable privacy protections for captured data, they share a fundamental limitation: they cannot prevent unauthorized sensing in the first place. An adversary can still collect the raw signals before any privacy protection is applied.
@@ 2.2.2. Real-Time Signal Protection
Jamming approaches inject noise or interference to degrade signal quality(decrease SNR) for unauthorized receivers. For audio privacy, ultrasonic jamming exploits microphone nonlinearity: high-frequency ultrasonic signals cause microphones to produce audible-range noise, corrupting recordings. Chen et al.(Chen et al., 2020) developed a wearable bracelet with 24 ultrasonic transducers providing omnidirectional microphone jamming. Yu et al.(Yu et al., 2025) improved efficiency through adaptive jamming that analyzes speech characteristics in real-time and generates time-frequency interference patterns matched to audio content. NUSGuard (Qiao et al., 2025) introduces temporal selectivity by detecting when users interact with authorized voice assistants, temporarily disabling jamming during those interactions. However, jamming approaches face fundamental limitations. Jamming is inherently modality-specific: it protects against one type of sensing signal, for example acoustic, but not RF or other sensing modalities, and vice versa, and we do not have control over what modality an unauthorized sensor uses.
@@
Obfuscation techniques inject plausible decoy information rather than noise, making it difficult to distinguish true signals from fake alternatives(Brunton and Nissenbaum, 2015). The core principle is to hide the real signal among believable decoys. For human tracking, RF-Protect (Shenoy et al., 2022) introduced a hardware reflector-based approach that injects phantom humans into device-free tracking systems. The system uses specially designed reflectors to modify radio waves and create reflections at arbitrary locations, combined with a generative model to create realistic human trajectories. For activity recognition, Argyriou (Argyriou, 2023) demonstrated that synthetic motion patterns can obfuscate human micro-Doppler signatures in passive radar. For vital sign protection, Anti-Sensing (Oshim et al., 2025) uses wearable oscillators that generate motion patterns mimicking natural cardiac motion, creating decoy signals that mislead radar-based heartbeat detection. While the approach successfully generates realistic oscillatory patterns, it blocks all radar sensors indiscriminately without providing selective access for authorized monitoring. VitalHide (Gao et al., 2025) recently explored whether phone vibrators and smart textile actuators could generate physical obfuscation for vital signs, creating decoy heartbeat signals to confuse unauthorized sensors while potentially allowing authorized devices to filter them out. The proof-of-concept showed that vibration-based obfuscation could reduce unauthorized detection accuracy. However, this work remained at the conceptual demonstration stage, lacking theoretical foundations, formal security analysis, and systematic validation of the selective access mechanism across different sensing modalities.
@@
Frequency Modulated Continuous Wave (FMCW) (Stove, 1992) has emerged as the preferred technology for contactless vital sign monitoring due to its ability to detect sub-millimeter movements with high precision without the need for sampling signal at carrier frequency. This capability enables FMCW radar to be implemented across a variety of signal modalities. In this section, we cover the FMCW radar theory and signal processing techniques used for extracting vital signs. We will then explain the design of PrivyWavethat can protect the subjects from unauthorized sensors that run these algorithms to extract physiological signals.
@@ 3.2. Signal Processing Methods for Vital Sign Extraction
Peak detection algorithms (Alizadeh et al., 2019) identify local maxima and minima in $x(t)$ corresponding to breathing cycles or cardiac phases, determining instantaneous rates from time intervals between peaks. FFT-based methods apply bandpass filtering to isolate specific frequency ranges (0.1–0.5 Hz for breathing, 0.8–2.0 Hz for heart rate), identifying peak frequencies as vital sign rates (Alizadeh et al., 2019; Lee et al., 2019). Time-frequency decomposition techniques  EMD/EEMD (Xu et al., 2022), VMD (Zhao et al., 2023), and wavelet transforms (Wang et al., 2023b) decompose the displacement signal into simpler signal components for analysis.
@@
Deep learning approaches train neural networks on the features that include the displacement $x(t)$ (Wang et al., 2023a; Kim et al., 2024) to estimate vital signs. These methods handle complex scenarios including multi-person environments (Wang et al., 2020) and low SNR conditions, but require substantial labeled data and may not generalize across deployment environments.
@@ 5.1. Experiment Setup
For unauthorized detection, we implement state-of-the-art heart rate measurement algorithms for mmWave (Alizadeh et al., 2019) and acoustic sensing (Qian et al., 2018). For authorized detection, we apply the same algorithms but first remove the known decoy frequencies using narrow band stop filters centered at $f_{1},\ldots,f_{p}$ as specified by the key $k$. Heart rate is computed from the displacement signal using the average RR interval over the 30-second recording period: $\text{Heart Rate (bpm)}=\frac{60}{\text{Average RR Interval}},$ and we evaluate performance using mean absolute error (MAE) between the measured heart rate and the Polar H10 ground truth.
@@ 5.2.3. Results
For mmWave Sensing: The unauthorized devices were effectively deceived. These sensors exhibited a high MAE of 21.3 $\pm$ 10.7 BPM, which confirms they consistently influenced by the decoy frequencies instead of the true heart rate. In contrast, the authorized device, which used the cryptographic key to filter out the decoy signals, achieved a low MAE of 5.8 $\pm$ 5.2 BPM. To put these errors into context, the reported heart-rate error of (Alizadeh et al., 2019) is $20\%$, while our errors converted into a percentage is $7.5\%$ for the authorized and $26\%$ for the unauthorized cases. So for mmWave, we are able to successfully preserve the utility of wireless heart-rate monitoring while providing obfuscation guaratee.
For Acoustic Sensing: The protection was even more robust, with unauthorized devices showing a very high MAE of 42.0 $\pm$ 12.4 BPM. The authorized device’s performance was slightly degraded compared to mmWave, with an MAE of 9.7 $\pm$ 7.3 BPM. The reported maximum heart-rate error of (Qian et al., 2018) is 3 BPM. This reduced accuracy for acoustic sensing (affecting both authorized and unauthorized measurements, as seen in the high error bars in Figure 5) is likely due to the physical properties of the modality. In our study, the users were just asked to sit in front of the system, but they did natural small movements while data collection, which also contributes to error while the reported method was more restricted during data collection. Acoustic signals have lower penetration through clothing and are more susceptible to environmental noise and multipath interference, resulting in a lower signal-to-noise ratio (SNR) compared to mmWave at the same range.
@@ 6.1. Motion Scenarios
Our evaluation focused on scenarios where participants remained stationary during measurements. This design choice reflects both the technical state of wireless heartbeat sensing and the most critical privacy threats in everyday life. Existing sensing algorithms achieve highest accuracy when subjects are stationary, as motion artifacts introduce significant noise that degrades detection (Parralejo et al., 2025). Also, these stationary conditions align with privacy-sensitive environments where individuals are most vulnerable to unauthorized monitoring: public transportation (buses, trains, airplanes), workplaces (offices, meeting rooms), public spaces (waiting rooms, restaurants, cafes, bars). In these settings, people remain relatively stationary for extended periods, often unaware of their surroundings, while potential adversaries have stable sensing conditions and ample time to collect high-quality physiological data. For instance, a malicious actor in a coffee shop could continuously monitor customers’ heart rates, or an unauthorized device in a shared office could track colleagues’ stress levels throughout the workday. By demonstrating effective protection in stationary settings, we address the harder and more prevalent threat case. In the future we will verify the effectiveness of these algorithms when the subject is moving.
@@ 6.2. Active Side-Channel Attacks
Our threat model assumes passive adversaries who observe wireless signals without actively probing the system. However, PrivyWave remains vulnerable to active side-channel attacks (Standaert, 2009) where an adversary could attempt to distinguish real vital signs from decoys through active stimulus-response probing. For example, an attacker might induce a sudden startle response and observe which signal components react, since the user’s physiological response would change while mechanical decoys would not. Defending against such active attacks represents a fundamental challenge that requires different approaches beyond physical obfuscation, such as detecting and responding to active probing attempts. Exploring defenses against active attacks is an potential direction for future work.

### v2

@@ 1. Introduction
Wireless sensing technologies have advanced rapidly in recent years. Systems using acoustic (nandakumar2015contactless), WiFi (abdelnasser2015ubibreathe), and millimeter-wave (mmWave) (yang2017vital) sensors can now monitor vital signs without any physical contact, detecting chest movements as small as few millimeters to measure breathing and heartbeat. These capabilities enable critical healthcare applications, such as detecting sleep apnea (nandakumar2015contactless), identifying opioid overdoses (nandakumar2019opioid), even detecting cardiac arrest and calling emergency services (chan2019contactless) all without requiring users to wear sensors.
However, these same technological capabilities that promise healthcare benefits also create privacy risks. Any ambient device such as a WiFi router, smart speaker, home assistant or someone’s smartphone, can potentially become a covert physiological signal monitor, silently extracting vital signs without user awareness or consent. This threat is particularly concerning because vital signs reveal far more than basic metrics, they could expose emotional states (zhao2016emotion), or stress levels (ha2021wistress) that may even indicate mental health conditions. In addition, some wireless signals are invisible and penetrate walls and clothing, creating an omnipresent surveillance risk which is hard for users to detect or avoid. So as wireless sensing becomes ubiquitous, developing solutions that maintain its practical benefits while preventing unauthorized surveillance is a necessity.
Existing privacy protection approaches for wireless signals fall into two categories: post-collection data processing and real-time protection. Post-collection approaches (shi2010prisense; hassan2019differential; liu2023application) protect data after acquisition through techniques such as data aggregation (shi2010prisense), differential privacy (hassan2019differential), and signal tampering (liu2023application). However, these methods cannot prevent unauthorized sensing in the first place, as adversaries can collect raw signals in real-time before any post-collection protection is applied. For real-time protection, jamming methods (chen2020wearable; yu2025dynamic; qiao2025nusguard) inject noise to degrade signal quality, but they are modality-specific—protecting against one type of sensor (e.g., acoustic) but not others (e.g., radio-frequency or RF). Anti-Sensing (oshim2025anti) uses wearable oscillators to mislead radar-based heartbeat detection, but blocks all sensors indiscriminately. VitalHide (gao2025vitalhide) presented the first approach for selective protection using vibration-based obfuscation. However, this work remained conceptual, lacking theoretical foundations, formal privacy analysis, and systematic validation.
@@ 2. Related Work
PriSense (shi2010prisense) Sensor Data N/A
Diff. Privacy (hassan2019differential) Sensor Data N/A
mmFilter (liu2023application) mmWave Radar N/A
@@
Wearable Jammer (chen2020wearable) Audio
Dynamic Jamming (yu2025dynamic) Audio
NUSGuard (qiao2025nusguard) Voice
@@
RF-Protect (shenoy2022rf) Human Tracking
Radar Obfus. (argyriou2023obfuscation) Activity
Anti-Sensing (oshim2025anti) Vital Signs
VitalHide (gao2025vitalhide) Vital Signs
@@ 2.1. Wireless Vital Sign Sensing and Privacy Implications
Wireless vital sign sensing has evolved into reliable systems with high accuracy. Many smart devices have gained the ability to accurately detect vital signs without physical contact. Nandakumar et al. (nandakumar2015contactless) demonstrated that smartphones can be turned into vital sign monitors by leveraging the acoustic sensors in the device. WiFi-based approaches extract breathing and heartbeat from the RF signals reflected by the subject (abdelnasser2015ubibreathe; liu2015tracking). The mmWave sensors, commonly deployed for presence detection, can also measure breathing and heartbeat with high precision (yang2017vital; gong2021rf).These sensing capabilities have enabled beneficial applications in everyday life. Researchers have developed systems for sleep apnea detection (nandakumar2015contactless), which can alert users to potentially dangerous breathing interruptions during sleep. Contactless vital sign monitoring has been proposed for overdose detection (nandakumar2019opioid), enabling earlier intervention in emergency situations. Cardiac arrest detection systems (chan2019contactless) can automatically alert emergency services when abnormal vital signs are detected. However, vital signals carry far more information than just physiological measurements. Stress levels can be inferred from vital signals (ha2021wistress), emotional states can be detected from combined respiratory and cardiac patterns (zhao2016emotion), and cognitive load during mental tasks can be assessed through cardiovascular responses (solhjoo2019heart). As sensing accuracy continues to improve, the precision of these inferences also increases, making the privacy risks more severe.
@@ 2.2.1. Post-Collection Data Processing
There are a set of algorithms that focuses on protecting data after it has been collected by sensors. These approaches assume that sensing has already occurred and apply privacy-preserving techniques during subsequent processing, transmission, or storage stages. PriSense (shi2010prisense) introduced privacy-preserving data aggregation for sensor networks using data slicing, where each sensor splits its reading into multiple shares distributed to randomly selected cover nodes, protecting individual data unless the aggregation server colludes with all cover nodes. Differential privacy approaches (hassan2019differential) provide stronger theoretical guarantees by adding calibrated noise to sensor data datasets, ensuring that the presence or absence of any individual’s data cannot be reliably determined while preserving aggregate statistics. For mmWave radar-based sensing, mmFilter (liu2023application) applies signal reversion techniques that tamper with radar data after collection but before transmission to sensing processors. While these post-collection methods provide valuable privacy protections for captured data, they share a fundamental limitation: they cannot prevent unauthorized sensing in the first place. An adversary can still collect the raw signals before any privacy protection is applied.
@@ 2.2.2. Real-Time Signal Protection
Jamming approaches inject noise or interference to degrade signal quality(decrease SNR) for unauthorized receivers. For audio privacy, ultrasonic jamming exploits microphone nonlinearity: high-frequency ultrasonic signals cause microphones to produce audible-range noise, corrupting recordings. Chen et al.(chen2020wearable) developed a wearable bracelet with 24 ultrasonic transducers providing omnidirectional microphone jamming. Yu et al.(yu2025dynamic) improved efficiency through adaptive jamming that analyzes speech characteristics in real-time and generates time-frequency interference patterns matched to audio content. NUSGuard (qiao2025nusguard) introduces temporal selectivity by detecting when users interact with authorized voice assistants, temporarily disabling jamming during those interactions. However, jamming approaches face fundamental limitations. Jamming is inherently modality-specific: it protects against one type of sensing signal, for example acoustic, but not RF or other sensing modalities, and vice versa, and we do not have control over what modality an unauthorized sensor uses.
@@
Obfuscation techniques inject plausible decoy information rather than noise, making it difficult to distinguish true signals from fake alternatives(brunton2015obfuscation). The core principle is to hide the real signal among believable decoys. For human tracking, RF-Protect (shenoy2022rf) introduced a hardware reflector-based approach that injects phantom humans into device-free tracking systems. The system uses specially designed reflectors to modify radio waves and create reflections at arbitrary locations, combined with a generative model to create realistic human trajectories. For activity recognition, Argyriou (argyriou2023obfuscation) demonstrated that synthetic motion patterns can obfuscate human micro-Doppler signatures in passive radar. For vital sign protection, Anti-Sensing (oshim2025anti) uses wearable oscillators that generate motion patterns mimicking natural cardiac motion, creating decoy signals that mislead radar-based heartbeat detection. While the approach successfully generates realistic oscillatory patterns, it blocks all radar sensors indiscriminately without providing selective access for authorized monitoring. VitalHide (gao2025vitalhide) recently explored whether phone vibrators and smart textile actuators could generate physical obfuscation for vital signs, creating decoy heartbeat signals to confuse unauthorized sensors while potentially allowing authorized devices to filter them out. The proof-of-concept showed that vibration-based obfuscation could reduce unauthorized detection accuracy. However, this work remained at the conceptual demonstration stage, lacking theoretical foundations, formal security analysis, and systematic validation of the selective access mechanism across different sensing modalities.
@@
Frequency Modulated Continuous Wave (FMCW) (stove1992linear) has emerged as the preferred technology for contactless vital sign monitoring due to its ability to detect sub-millimeter movements with high precision without the need for sampling signal at carrier frequency. This capability enables FMCW radar to be implemented across a variety of signal modalities. In this section, we cover the FMCW radar theory and signal processing techniques used for extracting vital signs. We will then explain the design of PrivyWavethat can protect the subjects from unauthorized sensors that run these algorithms to extract physiological signals.
@@ 3.2. Signal Processing Methods for Vital Sign Extraction
Peak detection algorithms (alizadeh2019remote) identify local maxima and minima in $x(t)$ corresponding to breathing cycles or cardiac phases, determining instantaneous rates from time intervals between peaks. FFT-based methods apply bandpass filtering to isolate specific frequency ranges (0.1–0.5 Hz for breathing, 0.8–2.0 Hz for heart rate), identifying peak frequencies as vital sign rates (alizadeh2019remote; lee2019novel). Time-frequency decomposition techniques  EMD/EEMD (xu2022non), VMD (zhao2023accurate), and wavelet transforms (wang2023slprof) decompose the displacement signal into simpler signal components for analysis.
@@
Deep learning approaches train neural networks on the features that include the displacement $x(t)$ (wang2023here; kim2024heartbeatnet) to estimate vital signs. These methods handle complex scenarios including multi-person environments (wang2020vimo) and low SNR conditions, but require substantial labeled data and may not generalize across deployment environments.
@@ 5.1. Experiment Setup
For unauthorized detection, we implement state-of-the-art heart rate measurement algorithms for mmWave (alizadeh2019remote) and acoustic sensing (qian2018acousticcardiogram). For authorized detection, we apply the same algorithms but first remove the known decoy frequencies using narrow band stop filters centered at $f_{1},\ldots,f_{p}$ as specified by the key $k$. Heart rate is computed from the displacement signal using the average RR interval over the 30-second recording period: $\text{Heart Rate (bpm)}=\frac{60}{\text{Average RR Interval}},$ and we evaluate performance using mean absolute error (MAE) between the measured heart rate and the Polar H10 ground truth.
@@ 5.2.3. Results
For mmWave Sensing: The unauthorized devices were effectively deceived. These sensors exhibited a high MAE of 21.3 $\pm$ 10.7 BPM, which confirms they consistently influenced by the decoy frequencies instead of the true heart rate. In contrast, the authorized device, which used the cryptographic key to filter out the decoy signals, achieved a low MAE of 5.8 $\pm$ 5.2 BPM. To put these errors into context, the reported heart-rate error of (alizadeh2019remote) is $20\%$, while our errors converted into a percentage is $7.5\%$ for the authorized and $26\%$ for the unauthorized cases. So for mmWave, we are able to successfully preserve the utility of wireless heart-rate monitoring while providing obfuscation guaratee.
For Acoustic Sensing: The protection was even more robust, with unauthorized devices showing a very high MAE of 42.0 $\pm$ 12.4 BPM. The authorized device’s performance was slightly degraded compared to mmWave, with an MAE of 9.7 $\pm$ 7.3 BPM. The reported maximum heart-rate error of (qian2018acousticcardiogram) is 3 BPM. This reduced accuracy for acoustic sensing (affecting both authorized and unauthorized measurements, as seen in the high error bars in Figure 5) is likely due to the physical properties of the modality. In our study, the users were just asked to sit in front of the system, but they did natural small movements while data collection, which also contributes to error while the reported method was more restricted during data collection. Acoustic signals have lower penetration through clothing and are more susceptible to environmental noise and multipath interference, resulting in a lower signal-to-noise ratio (SNR) compared to mmWave at the same range.
@@ 6.1. Motion Scenarios
Our evaluation focused on scenarios where participants remained stationary during measurements. This design choice reflects both the technical state of wireless heartbeat sensing and the most critical privacy threats in everyday life. Existing sensing algorithms achieve highest accuracy when subjects are stationary, as motion artifacts introduce significant noise that degrades detection (parralejo2025challenges). Also, these stationary conditions align with privacy-sensitive environments where individuals are most vulnerable to unauthorized monitoring: public transportation (buses, trains, airplanes), workplaces (offices, meeting rooms), public spaces (waiting rooms, restaurants, cafes, bars). In these settings, people remain relatively stationary for extended periods, often unaware of their surroundings, while potential adversaries have stable sensing conditions and ample time to collect high-quality physiological data. For instance, a malicious actor in a coffee shop could continuously monitor customers’ heart rates, or an unauthorized device in a shared office could track colleagues’ stress levels throughout the workday. By demonstrating effective protection in stationary settings, we address the harder and more prevalent threat case. In the future we will verify the effectiveness of these algorithms when the subject is moving.
@@ 6.2. Active Side-Channel Attacks
Our threat model assumes passive adversaries who observe wireless signals without actively probing the system. However, PrivyWave remains vulnerable to active side-channel attacks (standaert2009introduction) where an adversary could attempt to distinguish real vital signs from decoys through active stimulus-response probing. For example, an attacker might induce a sudden startle response and observe which signal components react, since the user’s physiological response would change while mechanical decoys would not. Defending against such active attacks represents a fundamental challenge that requires different approaches beyond physical obfuscation, such as detecting and responding to active probing attempts. Exploring defenses against active attacks is an potential direction for future work.